# Bounds on $k$-hash distances and rates of linear codes

### v2

@@ Abstract
In this paper, we bound the rate of linear codes in ${\mathbb{F}}_{q}^{n}$ with the property that any $k\leq q$ codewords are all simultaneously distinct in at least $d_{k}$ coordinates. For the particular case $d_{k}=1$, this leads to bounds on the rate of linear $q$-ary $k$-hash codes which generalize, with a simpler proof, results recently obtained for the case $q=k=3$ by Pohoata and Zakharov and by Bishnoi, D’haeseleer, Gijswijt and Potukuchi. We finally discuss some related open problems on the list-decoding zero-error capacity of discrete memoryless channels.
@@ I Introduction
Let $q\geq k\geq 3$ and $n\geq 1$ be integers. The notion of minimum $k$-hash distance of a $q$-ary code $C$ of length $n$ was introduced in [1]. In particular, given $k$ codewords, let their joint $k$-hash distance be defined as the number of coordinates in which they are all distinct. We say that a code $C$ has $k$-hash distance $d_{k}$ if $d_{k}$ is minimum $k$-hash distance among all $k$-tuples of distinct codewords. We also define $\delta_{k}:=d_{k}/n$ as the relative $k$-th hash distance of the code. Clearly, $d_{2}$ is the usual minimal Hamming distance.
In what follows, for every positive integers $b\leq a$, we define $a^{\underline{b}}=a(a-1)\cdots(a-b+1)$ and for every real $\alpha$ we define $(\alpha)^{+}=\max(0,\alpha)$.
In [1], the authors provide the following bounds for codes with $k$-hash distance $d_{k}$.
Theorem 1 ([1])
Let $C$ be a $q$-ary code with $k$-th hash distance $d_{k}$. Then
$|C|\leq(k-1)q^{\left\lceil\frac{n-d_{k}+1}{k-1}\right\rceil}\,.$
Theorem 2 ([1])
Let $C$ be a $q$-ary code with $k$-th hash distance $d_{k}$. Then
$|C|\leq d_{k}\binom{k}{2}\left(\frac{q}{k-1}\right)^{n-\frac{q^{k}}{q^{\underline{k}}}(d_{k}-1)}.$
It can be seen that for $q\gg k$ and $n$ large enough, Theorem 1 provides a better bound than Theorem 2.
A case of special importance is that of codes with minimum $k$-hash distance $d_{k}\geq 1$, usually simply called $(q,k)$-hash codes. Most of the literature on $k$-hashing is devoted to this particular case. The problem of finding upper and lower bounds for the maximum size of $(q,k)$-hash codes is a fundamental problem in theoretical computer science and information theory. It appears, as the name suggests, in the study of families of perfect hash functions and in the study of the zero-error capacity of some discrete channels with list decoding, see [2, 3, 4, 5, 6] for more details (see also [7] for a related problem). We thus devote some attention to this case and will return later on to the general problem of possibly larger values of $d_{k}$.
An elementary double counting argument, as shown in [3], gives the following bound on the cardinality of $(q,k)$-hash codes:
@@
It can be seen that this bound improves the one of Theorem 2 for $d_{k}=1$. In 1984 Fredman and Komlós [2] improved the bound in (1) for every $q=k\geq 4$ and sufficiently large $n$, obtaining the following result:
@@
Fredman and Komlós also provided, using standard probabilistic methods, the following lower bound:
@@
A generalization of the upper bounds given in equations (1) and (2) was derived by Körner and Marton [3] in the form
@@
In 1998, Blackburn and Wild [5] (see also [1]) improved the bound of Körner and Marton for every $q$ sufficiently larger than $k$, proving
@@
This bound matches the one given in Theorem 1 for $d_{k}=1$.
Much effort has been spent during the years to refine the bounds given in (1), (2) and (4). Exponential improvements were obtained in [8, 4, 9] for the case of $q=k=4$, in [10] for $q=k=5,6$, and in [3, 11, 12, 6] for $q\geq k\geq 5$. However, to the best of our knowledge, for $q$ sufficiently larger than $k$ no improvements over the upper bound (5) have been obtained. Improvements on the lower bound given in (3) on the largest size of $(q,k)$-hash codes have been recently obtained in [13] for $q\in[4,15]$, all integers between $17$ and $25$, and for a sufficiently large $q$. For $q=k=3$ the best known lower bound is due to Körner and Marton [3] and it is equal to $(9/5)^{n/4+o(1)}$.
In contrast, no exponential improvement has been made on the simple bound given in (1) for the so called trifference problem, that is, $q=k=3$. Until recently, only improvements on the multiplicative constant had been obtained (see [14, 15]), while a polynomial improvement has been obtained in a beautiful recent work by Bhandhari and Kheta [16].
However, if we restrict the codes to be linear, i.e., we require $C$ to be a linear subspace of ${\mathbb{F}}_{q}^{n}$, exponential improvements have been first obtained in [1] where the authors provide the following recursive bound.
Theorem 3 ([1])
$d_{s+1}\leq\left(d_{s}-m+1\right)^{+}\,.$
Iterating Theorem 3 we get
$d_{k}\leq\left(d_{2}-(k-2)(m-1)\right)^{+}\,.$ (6)
Whenever $m/n\to R$ and $d_{k}/n\to\delta_{k}>0$ as $n\to\infty$, the previous equation gives the following result.
Theorem 4
Let $C$ be a linear code in ${\mathbb{F}}_{q}^{n}$ of rate $R$ and minimum relative $k$-th hash distance $\delta_{k}>0$ and minimum Hamming relative distance $\delta$. Then
$R\leq\frac{\delta}{k-2}+O\left(\frac{1}{n}\right)\,.$
Due to Theorem 4 for $q=k=3$ and the linear programming bound for $3$-ary codes by Aaltonen [17] (see eq. (11) later on), we can derive the following upper bound on the cardinality of linear trifferent codes.
$|C|$ $\leq 3^{n/2.8272+o(n)}\approx 1.4749^{n}\,.$
Improved upper bounds on the rate of linear trifferent codes have been recently considered in [18], where it was proved that for some $\epsilon>0$,
@@
where the constant $1.2731$ is the numerical solution of an equation that we will explain in Section II-B. The authors also showed that there exist linear trifferent codes of length $n$ and size $\frac{1}{3}(9/5)^{n/4}$, matching, asymptotically in $n$, the best known lower bound on trifferent codes (without the linearity constraint) obtained in [3].
@@
We note that when $q$ is small compared to $k>3$, no linear $k$-hash codes of dimension $2$ exist. In particular, Blackburn and Wild [5] showed that this is true for every $q\leq 2k-4$ (so, in particular, the case $q=k>3$ is of no interest). The authors in [20] improved this result for $k\geq 9$ showing that when $q$ is a square, $q>25$, no linear $k$-hash codes of dimension $2$ exist whenever $q\leq\left(\frac{k-1}{2}\right)^{2}$. Hence in these regimes, we know that linear $k$-hash codes in ${\mathbb{F}}_{q}^{n}$ are relatively simple objects, since their asymptotic rates are equal to zero.
@@
In this paper we provide upper bounds on the rate of linear codes in ${\mathbb{F}}_{q}^{n}$ with minimum relative asymptotic distance $\delta_{k}>0$, for general values of $q\geq k\geq 3$. The method can be stretched to consider $d_{k}\geq 1$, thus also providing upper bounds on $q$-ary linear $k$-hash codes. For $q\geq k>3$ these are the first known (non-trivial) such bounds. For $q=k=3$ they recover the best known result of [19] given in equation (8), but with a simpler proof. Also, in the range of $q$ much larger than $k$, they improve the general bound of equation (5) (in terms of code rate as $n\to\infty$).
@@
The paper is structured as follows. In Section II we review the case of $q=k=3$ which is of special importance for comparison with the recent literature, showing how our method applies in this case to recover the known results in a simpler way. In Section III we present the general method for $q\geq k\geq 3$ and $d_{k}\geq 1$. In Section IV we discuss the derived bounds and compare them with those in the literature. In Section V we discuss some connections of the proposed method with some open problems on list decoding.
@@ II The case $q=k=3$
The case $q=k=3$ deserves special attention being, as mentioned in the introduction, the simplest yet the most difficult case to study. In this Section, we present a re-derivation of the best known results in the literature. We first present a random-coding proof that the achievability bound of Körner and Marton [3] also holds for linear codes, proof which differs from the one given in [19]. Then we present our proof of the converses in [18] and [19]. The reason is for rederiving these converses is two-fold. First, we believe the proposed proof is conceptually simpler, and more in the spirit of a traditional information-theoretic toolbox. Second, the method used to derive the upper bound extends to the general case $q\geq k\geq 3$, contrarily to the one previously used in the literature. This is indeed the contribution of the rest of the paper.
@@ II-A Lower Bound
The best known lower bound on the rate of general (not necessarily linear) codes is due to Körner and Marton [3] and says that
$R\geq\frac{1}{4}\log\frac{9}{5}-o(1)\,.$
For the sake of completeness, we review their result, which uses a standard expurgation trick (see also Elias [21]).
Consider the so-called tetracode [^1] See also [21]. In [3] a different but equivalent code is used., a trifferent code of length $4$ with the following $9$ codewords $b_{0},\ldots,b_{8}$
$\begin{array}[]{cccc}0&0&0&0\\ 0&1&2&1\\ 0&2&1&2\\ 1&0&2&2\\ 1&1&1&0\\ 1&2&0&1\\ 2&0&1&1\\ 2&1&0&2\\ 2&2&2&0\\ \end{array}$ (9)
For $n$ a multiple of 4, consider any $9$-ary code $C^{\prime}$ of length $n/4$ over the alphabet $\{0,1,\ldots,8\}$. Let then $C$ be a ternary code of length $n$ obtained by replacing each entry of value $i$ in $C^{\prime}$ with the corresponding cordeword $b_{i}$, $i\in\{0,1,,\ldots,8\}$. Note that, since the tetracode is trifferent, $C$ is a $(3,3)$-hash code if $C^{\prime}$ is a $(9,3)$-hash code. Let now $C^{\prime\prime}$ be a random $9$-ary code of length $n/4$ with $2M$ independent codewords composed of uniform i.i.d. symbols. The probability that any three codewords in $C^{\prime\prime}$ are not $3$-hashed is precisely
@@
The above argument shows that the rate $R=\frac{1}{4}\log(9/5)$ is achievable by means of general codes. It was proved in [19] that this rate can also be achieved by linear codes [^2] We neglect the coefficients in the bounds for $|C|$ here. In [19] the authors prove $|C|\geq\frac{1}{3}(9/5)^{n/4}$. While this work was already in preparation, the authors added in the revised version [19] that, according to a referee, the bound $|C|\geq(9/5)^{n/4}$ for linear codes can be proved along the lines of Körner and Marton. We could not verify this stronger statement.. Interestingly enough, the proof used in [19] uses a different procedure which goes through upper bounds on affine $2$-blocking sets. In particular, no use is made of the tetracode. It is rather interesting that the same asymptotic result is obtained with that approach.
Here, we show that the achievability of the rate $R=\frac{1}{4}\log(9/5)$ for linear codes can also be proved using the Körner-Marton method based on the tetracode. First note that the tetracode in (9) is linear as a code in ${\mathbb{F}}_{3}$ with generator matrix
@@
So, any codeword $b_{i}$ can be represented uniquely by the information vector $a_{i}\in{\mathbb{F}}_{3}^{2}$ such that $b_{i}=a_{i}\,G_{T}$.
Consider now a random linear code $C^{\prime}$ over ${\mathbb{F}}_{9}$ of length $n/4$, whose $k\times n/4$ generator matrix has uniform i.i.d. entries. The probability that the code is not a $(9,3)$-hash code equals the probability that there are at least two distinct non-zero vectors $u_{1}$ and $u_{2}$ in ${\mathbb{F}}_{9}^{k}$ such that the triplet $\{0,u_{1}G,u_{2}G\}$ is not $3$-hashed. If $g_{i}$ is the $i$-th column of $G$, then this happens if
@@
If $u_{1}$ and $u_{2}$ are linearly dependent, then |$|\{0,u_{1}g_{i},u_{2}g_{i}\}|\leq 2$ if and only if $u_{1}g_{i}=0$, which happens with probability $1/9$. If instead $u_{1}$ and $u_{2}$ are linearly independent, then $u_{1}g_{i}$ and $u_{2}g_{i}$ take on any pair of values with equal probability, and so (|$\mathbb{P}(|\{0,u_{1}g_{i},u_{2}g_{i}\}|\leq 2)=25/81$. So, the probability that $\{0,u_{1}G,u_{2}G\}$ is not $3$-hashed is less than $\left(\frac{25}{81}\right)^{n/4}=P$ (as before). Hence, the expected number of such bad $(u_{1},u_{2})$ pairs is less than $\binom{9^{k}}{2}P<9^{2k}P/2$. This will be less than one if
$9^{k}\leq\sqrt{2}\left(\frac{9}{5}\right)^{n/4}\,.$
@@
In that case, there exists at least one linear $(9,3)$-hash code $C^{\prime}$ with $M=9^{k}$ codewords. This code can then be turned into a linear trifferent code of lentgh $n$ (over ${\mathbb{F}}_{3}$) by interpreting the codeword symbols as elements in ${\mathbb{F}}_{3}^{2}$ and multiplying them by the generator matrix of the tetracode $G_{T}$. This shows that the rate $R=\frac{1}{4}\log(9/5)$ can also be achieved by linear trifferent codes.

### v3

@@ Abstract
In this paper, we bound the rate of linear codes in ${\mathbb{F}}_{q}^{n}$ with the property that any $k\leq q$ codewords are all simultaneously distinct in at least $d_{k}$ coordinates. For the case of particular interest $q=k=3$ we recover, with a simpler proof, state of the art results in the case $d_{3}=1$ and new bounds for $d_{3}>1$. We finally discuss some related open problems on the list-decoding zero-error capacity of discrete memoryless channels.
@@ I Introduction
In general, in the whole paper, when the base of a logarithm is not specified it is assumed to be the cardinality $q$ of the code alphabet being considered. Also, in what follows, for every positive integers $b\leq a$, we use the notation $a^{\underline{b}}$ for the falling factorial $a^{\underline{b}}=a(a-1)\cdots(a-b+1)$ and for every real $\alpha$ we write the positive part as $(\alpha)^{+}=\max(0,\alpha)$.
Given $k\leq q$ fixed codewords of a $q$-ary code, we say that they are $k$-hashed if there is a coordinate where they are all pairwise distinct. A $q$-ary code is said to be a $(q,k)$-hash code if any $k$ distinct codewords are $k$-hashed. The problem of finding upper and lower bounds for the maximum size of $(q,k)$-hash codes is a fundamental problem in theoretical computer science and information theory. It appears, as the name suggests, in the study of families of perfect hash functions and in the study of the zero-error capacity of some discrete channels with list decoding, see [1, 2, 3, 4, 5] for more details (see also [6] for a related problem). We review some of the most important results in this direction, as it will be useful to give a context to the results presented in this paper.
An elementary double counting argument, as shown in [2], proves that the cardinality of $(q,k)$-hash codes is bounded as
@@
In 1984 Fredman and Komlós [1] improved the bound in (1) for every $q=k\geq 4$ and sufficiently large $n$, showing that
@@
They also provided, using standard probabilistic methods, an achievability result proving the existence of codes with cardinality lower bounded as
@@
Körner and Marton [2] presented a generalization of the upper bounds given in inequalities (1) and (2) which takes the form
@@
Note in particular that (2) corresponds to only using j=k2j=k-2 in the argument of the minimization in (4). For values of $q$ sufficiently larger than $k$, these bounds become weak; in 1998 Blackburn and Wild [4] (see also [7]) showed that they can be improved by the following bound
@@
[…]
$|C|\geq(9/5)^{n/4+o(1)}\,.$ (6)
In contrast, no exponential improvement has been made on the simple bound given in (1) for this case. Until recently, only improvements on the multiplicative constant had been obtained (see [14, 15]), while a polynomial improvement of the form
$|C|\leq\frac{c}{n^{2/5}}\left(\frac{3}{2}\right)^{n}$ (7)
for some constant $c$ has been obtained in a beautiful recent work by Bhandhari and Khetan [16].
Additional results were obtained for the case of linear codes, i.e., we require $C$ to be a linear subspace of ${\mathbb{F}}_{q}^{n}$. Improved upper bounds on the rate of linear trifferent codes have been recently considered in [17], where it was proved that for some $\epsilon>0$,
@@
where the constant $1.2731$ is the numerical solution of an equation that we will explain in Section II-B. The authors also showed that there exist linear trifferent codes of length $n$ and size $\frac{1}{3}(9/5)^{n/4}$, matching, asymptotically in $n$, the best known lower bound on trifferent codes (without the linearity constraint) obtained in [2]. We note that when $q$ is small compared to $k>3$, no linear $q$-ary $k$-hash codes of dimension $2$ exist. In particular, Blackburn and Wild [4] showed that this is true for every $q\leq 2k-4$ (so, in particular, the case $q=k>3$ is of no interest). The authors in [19] improved this result for $k\geq 9$ showing that when $q$ is a square, $q>25$, no linear $k$-hash codes of dimension $2$ exist whenever $q\leq\left(\frac{k-1}{2}\right)^{2}$. Hence in these regimes, we know that linear $k$-hash codes in ${\mathbb{F}}_{q}^{n}$ are relatively simple objects, since their asymptotic rates are equal to zero.
In this paper we focus on linear codes which satisfy a generalization of the $k$-hashing property, namely a notion of $k$-hash distance. The notion of minimum $k$-hash distance of a $q$-ary code $C$ of length $n$ was introduced in [7]. In particular, given $k$ codewords, let their joint $k$-hash distance be defined as the number of coordinates in which they are all distinct. We say that a code $C$ has $k$-hash distance $d_{k}$ if $d_{k}$ is the minimum $k$-hash distance among all $k$-tuples of distinct codewords. We also define $\delta_{k}:=d_{k}/n$ as the relative $k$-hash distance of the code. Clearly, $d_{2}$ is the usual minimal Hamming distance. In [7], the authors provide the following two upper bounds on the cardinality of (general) $q$-ary codes with $k$-hash distance $d_{k}$
$|C|$ $\leq(k-1)q^{\left\lceil\frac{n-d_{k}+1}{k-1}\right\rceil}\,$, (10)
$|C|$ $\leq d_{k}\binom{k}{2}\left(\frac{q}{k-1}\right)^{n-\frac{q^{k}}{q^{\underline{k}}}(d_{k}-1)}.$ (11)
Note that for $d_{k}=1$ equation (10) gives (5), while (11) recovers at the exponential order the bound in equation (1). More generally, for fixed $d_{k}$ and $q\gg k$, in equation (11) we have ${q^{k}}/{q^{\underline{k}}}\approx 1$, so that the bound grows for large $n$ as $(q/(k-1))^{n-d_{k}}$, while the bound in equation (10) has instead order of magnitude $q^{(n-d_{k})/(k-1)}$, thus being tighter in this regime. On the other hand, for fixed $q$ and $k$, the bound in (11) shows that positive rates are only possible for asymptotic relative distance $\delta_{k}\leq\frac{q^{\underline{k}}}{q^{k}}<1$. More recently, new bounds were derived in [20] independently and essentially at the same time of the first version of the present work (see some comparison in Section II-B).
In [7], specific results are also given for linear codes. In particular, it is shown that $q$-ary linear codes of length $n$ and dimension $m$ with minimum $k$-hash distance $d_{k-1}>0$ have $k$-hash distance $d_{k}$ bounded by
$d_{k}\leq\left(d_{k-1}-m+1\right)^{+}\,.$ (12)
Iterating equation (12) one gets
$d_{k}\leq\left(d_{2}-(k-2)(m-1)\right)^{+}\,.$ (13)
In the regime $m/n\to R$ as $n\to\infty$ with $d_{k}=1$, the previous equation shows that linear codes in ${\mathbb{F}}_{q}^{n}$ with minimum Hamming asymptotic relative distance $\delta_{2}$ have rate bounded as
$R\leq\frac{\delta_{2}}{k-2}+O\left(\frac{1}{n}\right)\,.$ (14)
Using (14) for $q=k=3$ combined with the linear programming bound for $3$-ary codes by Aaltonen [21] (see Section II-B for details), one can derive the following upper bound on the cardinality of linear trifferent codes.
$|C|$ $\leq 3^{n/2.8272+o(n)}\approx 1.4749^{n}\,.$
Note that this is already better than the best bound for general codes (7) but weaker than (8).
@@
In this paper we provide upper bounds on the rate of linear codes in ${\mathbb{F}}_{q}^{n}$ with minimum distance $d_{k}>0$, for general values of $q\geq k\geq 3$, where $d_{k}$ is either constant or grows linearly with $n$ so that $d_{k}/n\to\delta_{k}$. In particular, for $d_{k}=1$ we derive upper bounds on $q$-ary linear $k$-hash codes. For $q\geq k>3$ these are the first known (non-trivial) such bounds, while for $q=k=3$ they recover the best known result of [18] given in equation (9), but with a simpler proof. Also, in the range of $q$ much larger than $k$, they improve the general bound of equation (5) in terms of code rate as $n\to\infty$.
@@
The paper is structured as follows. In Section II we review the case of $q=k=3$ which is of special importance for comparison with the recent literature, showing how our method applies in this case. In Section III we present the general method for $q\geq k\geq 3$ and we discuss the derived bounds comparing them with those in the literature. In Section IV we discuss some connections of the proposed method with some open problems on list decoding.
@@ II The case $q=k=3$
The case $q=k=3$ deserves special attention being, as mentioned in the introduction, the simplest yet the most difficult case to study. In this Section, we present a re-derivation of the best known results in the literature for the case $d_{3}=1$, and we show how our proofs extend easily to give good results for the case $d_{3}>1$. In particular, we first present a random-coding proof that the achievability bound (6) of Körner and Marton [2] also holds for linear codes. This proof differs from the one given in [18], and we show how to extend it to also cover the case $d_{3}>1$. Then we present our proof, which we consider to be conceptually simpler and more in the spirit of a traditional information-theoretic toolbox, of the converses (8) and (9) of [17] and [18], showing how the method can be also used to cover the case $d_{3}>1$.
@@ II-A Lower Bound
The best known lower bound on the rate of general (not necessarily linear) trifferent codes remains to date Körner and Marton’s bound (6) derived in [2]. For the sake of completeness, we review their result, which uses a standard expurgation trick (see also Elias [22]). Consider the so-called tetracode [^1] See also [22]. In [2] a different but equivalent code is used., a trifferent code of length $4$ with the following $9$ codewords $b_{0},\ldots,b_{8}$
[…]
For $n$ a multiple of 4, consider any $9$-ary code $C^{\prime}$ of length $n/4$ over the alphabet $\{0,1,\ldots,8\}$. Let then $C$ be a ternary code of length $n$ obtained by replacing each entry of value $i$ in $C^{\prime}$ with the corresponding codeword $b_{i}$, $i\in\{0,1,,\ldots,8\}$. Note that, since the tetracode is trifferent, $C$ is a $(3,3)$-hash code if $C^{\prime}$ is a $(9,3)$-hash code. Let now $C^{\prime\prime}$ be a random $9$-ary code of length $n/4$ with $2M$ independent codewords composed of uniform i.i.d. symbols. The probability that any three codewords in $C^{\prime\prime}$ are not $3$-hashed is precisely
@@
The above argument shows that the rate $R=\frac{1}{4}\log(9/5)$ is achievable by means of general codes. It was proved in [18] that this rate can also be achieved by linear codes [^2] We are only interested in bounds on the rate of codes here, and not on subexponential coefficients in bounds for $|C|$. In [18] the authors prove $|C|\geq\frac{1}{3}(9/5)^{n/4}$. While this work was already in preparation, the authors added in the revised version [18] that, according to a referee, the bound $|C|\geq(9/5)^{n/4}$ for linear codes can be proved along the lines of Körner and Marton. We could not verify this stronger statement; because $|C|$ must be a power of $9$, our bound only proves the existence of a linear code with $|C|\geq\frac{\sqrt{2}}{9}(9/5)^{n/4}$.. Interestingly enough, the proof in [18] uses a different procedure which goes through upper bounds on affine $2$-blocking sets. In particular, no use is made of the tetracode. It is rather interesting that the same asymptotic result is obtained with that approach.
Here, we show that the achievability of the rate $R=\frac{1}{4}\log(9/5)$ for linear codes can also be proved using the Körner-Marton method based on the tetracode. First note that the tetracode in (15) is linear as a code over ${\mathbb{F}}_{3}$ with generator matrix
@@
So, any codeword $b_{i}$ can be represented uniquely by an information vector $a_{i}\in{\mathbb{F}}_{3}^{2}$ such that $b_{i}=a_{i}\,G_{T}$.
Consider now a random linear code $C^{\prime}$ over ${\mathbb{F}}_{9}$ of length $n/4$, whose $m\times n/4$ generator matrix has uniform i.i.d. entries. The probability that the code is not a $(9,3)$-hash code equals the probability that there are at least two distinct non-zero vectors $u_{1}$ and $u_{2}$ in ${\mathbb{F}}_{9}^{m}$ such that the triplet $\{0,u_{1}G,u_{2}G\}$ is not $3$-hashed. If $g_{i}$ is the $i$-th column of $G$, then this happens if
@@
If $u_{1}$ and $u_{2}$ are linearly dependent, then $|\{0,u_{1}g_{i},u_{2}g_{i}\}|\leq 2$ if and only if $u_{1}g_{i}=0$, which happens with probability $1/9$. If instead $u_{1}$ and $u_{2}$ are linearly independent, then $u_{1}g_{i}$ and $u_{2}g_{i}$ take on any pair of values with equal probability, and so $\mathbb{P}(|\{0,u_{1}g_{i},u_{2}g_{i}\}|\leq 2)=25/81$. So, the probability that $\{0,u_{1}G,u_{2}G\}$ is not $3$-hashed is at most $\left(\frac{25}{81}\right)^{n/4}=P$ (as before). Hence, the expected number of such bad $(u_{1},u_{2})$ pairs is less than $\binom{9^{m}}{2}P<9^{2m}P/2$. This will be less than one if
$9^{m}$ $\leq\sqrt{2}\,P^{-1/2}$
$=\sqrt{2}\left(\frac{9}{5}\right)^{n/4}\,.$
@@
In that case, there exists at least one linear $(9,3)$-hash code $C^{\prime}$ with $M=9^{m}$ codewords. This code can then be turned into a linear trifferent code of length $n$ (over ${\mathbb{F}}_{3}$) by interpreting the codeword symbols as elements in ${\mathbb{F}}_{3}^{2}$ and multiplying them by the generator matrix of the tetracode $G_{T}$. This shows that the rate $R=\frac{1}{4}\log(9/5)$ can also be achieved by linear trifferent codes.
The bound derived above holds for trifferent codes, that is codes with $3$-hash distance $d_{3}$ at least $1$. However, the method can be extended to bound the rate of codes with minimum distance at least $d_{3}>1$. We consider here the case where $d_{3}$ grows linearly with $n$ so that $d_{3}/n\to\delta_{3}>0$. We have the following result.
Theorem 1
For any $\delta_{3}>0$, ternary codes with asymptotic relative $3$-hash distance $\delta_{3}$ exist for all rates $R$ satisfying
$R<\frac{1}{8}D(p^{*}\|p)\,$,
where $D(\cdot\|\cdot)$ is the Kullback-Leibler divergence [23] (with $\log$ in base $3$), $p$ is the probability vector $p=(\frac{25}{81},\frac{48}{81},0,\frac{8}{81},0)$ and $p^{*}$ is defined by
$p_{j}^{*}=\frac{p_{j}3^{\alpha\cdot j}}{\sum_{h}p_{h}3^{\alpha\cdot h}}\,,\qquad j=0,1,2,3,4\,$,
with $\alpha$ such that $\sum_{j}j\cdot p_{j}^{*}=4\delta_{3}$.
Proof:
We just need to modify the argument presented for $d_{3}=1$ by considering the probability that any triplet of codewords $\{0,u_{1}G,u_{2}G\}$ in $\mathbb{F}_{9}^{n/4}$ correspond, after multiplication of each element by the tetracode matrix $G_{T}$, to ternary codewords with $3$-hash distance less then $\delta_{3}n$. This can be done by carefully checking the number of trifferences which correspond to different pairs of tetracode codewords.
Consider again a coordinate $i$ for the $9$-ary code $C^{\prime}$, and the pair of values $u_{1}g_{i}$ and $u_{2}g_{i}$, where $g_{i}$ is the $i$-th random column of $G$. Call $T_{i}$ the random variable which measures the number of coordinates where the two tetracode codewords corresponding to $u_{1}g_{i}$ and $u_{2}g_{i}$ differ and are both non-zero. Then, the analysis done for $d_{3}=1$ can be modified by replacing the probability $P$ used above with P’= P [∑_i=1^n/4 T_i ¡ d_3] , showing existence of linear codes with $3$-hash distance $\delta_{3}n$ for all rates $R<-\frac{1}{n}\log\sqrt{P^{\prime}}-o(1)$. To bound $P^{\prime}$ we only need to study the distribution of $T_{i}$. If $u_{1}$ and $u_{2}$ are linearly independent, then $u_{1}g_{i}$ and $u_{2}g_{i}$ take on all positive pairs with equal probability and one can check that $T_{i}$ is distributed over $\{0,1,2,3,4\}$ with distribution p=(2581, 4881, 0, 881, 0) . We can bound the probability $P^{\prime}$ to the first order in the exponent using Sanov’s theorem [23] whenever $\delta_{3}\leq E[T_{i}]/4=2/9$. We then find
$P^{\prime}\leq 3^{-\frac{n}{4}D(p^{*}\|p)+o(n)}\,$, (16)
where $D(\cdot\|\cdot)$ is the Kullback-Leibler divergence (with $\log$ in base $3$) and p_j^*=pj3α⋅jhph3α⋅h ,  j=0,1,2,3,4 , with $\alpha$ such that $\sum_{j}j\cdot p_{j}^{*}=4\delta_{3}$.
If $u_{1}$ and $u_{2}$ are linearly dependent, what mentioned for the case $d_{3}=1$ says that $T_{i}\geq 1$ with probability at least $8/9$. So, the probability $P^{\prime}$ is upper bounded by the Chernoff bound as $P^{\prime}\leq 3^{-\frac{n}{4}D(p^{\prime}\|p)}$, with $p^{\prime}=(4\delta_{3},1-4\delta_{3})$ and $p=(8/9,1/9)$. Numerical evaluations show that for all values of $\delta_{3}$ this upper bound is much smaller than the one in (16), which thus holds in general. Using $R<-\frac{1}{n}\log\sqrt{P^{\prime}}-o(1)$ with $P^{\prime}$ as in (16) we obtain the claimed result. ∎
For $\delta_{3}\to 0$, Theorem 1 recovers the achievability of Körner and Marton [2], and in general improves for all $\delta_{3}\in[0,2/9]$ the bound
$R<\frac{1}{2}D(\delta_{3}\|2/9)\,$
obtained with the same method but using random linear codes built directly from a ternary random $m\times n$ generator matrix with i.i.d. uniform entries [7, Theorem 2] (see Figure 1). Note that both bounds vanish at $\delta_{3}=2/9$, which is tight as proved by [7, Th. 5] and [20, Th. 1.7].
Figure 1: Achievable rates for linear ternary codes with $3$-hash distance $\delta_{3}$.